A top spins for 10.51 min, beginning with an angul
Answer #1

t = 10.51 min = 10.51*60 = 630.6 sec

wi = 10.31 rev/sec = 64.78 rad/sec

wf = 0

wf = wi + alpha*t

alpha = (wf - wi)/t

alpha = ?

alpha = (0 - 64.78)/630.6 = -0.1027 rad/sec^2

theta = ?

theta = wi*t + 0.5*alpha*t^2

theta = 64.78*630.6 - 0.5*0.1027*630.6^2

theta = 20430.61 rad

theta = 1170587.73 deg